# Quantum amplification and simulation of strong and ultrastrong coupling of light and matter

### v1

@@ 2.2.2 Collective amplification
In Sec. 2.2.1, we showed that a detuned parametric driving of a cavity enables the coupling between an atom and the squeezed cavity mode to be enhanced exponentially. Because exponentially many photons are converted into a single-photon state of the squeezed cavity mode, this parametric enhancement of the atom-field coupling can thus be understood as originating from the coupling of a single atom to many photons. In this section, we introduce an opposite enhancement mechanism, which is based on the coupling of a single photon to an ensemble containing many atoms or spins. To proceed, let us assume that the ensemble contains $N$ identical two-level atoms, as illustrated in Fig. 5. The coupling between the ensemble and a cavity mode is described by the interaction Hamiltonian
@@
Let us now apply the Holstein-Primakoff transformation Holstein and Primakoff (1940)
$S_{-}=\;\sqrt{N-s^{\dagger}s}s\quad{\rm and}\quad S_{+}=\;s^{\dagger}\sqrt{N-s^{\dagger}s},$ (69)
where $s$ and $s^{\dagger}$ are the bosonic annihilation and creation operators, which satisfy the commutation relation $\left[s,s^{\dagger}\right]=1$. In the low-excitation regime, where the average number of excited atoms is much smaller than the total number of atoms (i.e., $\langle s^{\dagger}s\rangle\ll N$), this transformation is further simplified to
$S_{-}\approx\;\sqrt{N}s\quad{\rm and}\quad S_{+}\approx\;\sqrt{N}s^{\dagger}.$ (70)
It is seen that the collective spin behaves as a quantum harmonic oscillator. In this case, the Hamiltonian $H_{\rm ens}^{\rm int}$ becomes
@@
The collective strong coupling between light and matter was first demonstrated experimentally with an ensemble of Rydberg atoms in Ref. Kaluzny et al. (1983), and has been a fundamental building block of quantum repeaters for long-distance quantum communication Duan et al. (2001); Kuzmich et al. (2003); van der Wal et al. (2003); Chou et al. (2005); Chanelière et al. (2005); Yuan et al. (2008); Hammerer et al. (2010); Sangouard et al. (2011); Grèzes et al. (2016); Jing et al. (2019); Yu et al. (2020); Pu et al. (2021). Moreover, a collective enhancement of light-matter interactions allows one to generate various nonclassical states in large ensembles, e.g., spin-squeezed states Kitagawa and Ueda (1993); Hald et al. (1999); Sørensen and Mølmer (2002); Leroux et al. (2010); Schleier-Smith et al. (2010); Ma et al. (2011); Dalla Torre et al. (2013); Tucker et al. (2020); Macrì et al. (2020); Qin et al. (2020); Groszkowski et al. (2020); Liu et al. (2023) and atomic Schrödinger cat states Agarwal et al. (1997); Massar and Polzik (2003); Qin et al. (2021). Thus, it also forms an essential ingredient of quantum metrology for high-precision measurements Pezzè et al. (2018).
@@
Recently, much attention has been focused on coupling nitrogen-vacancy (NV) electronic-spin ensembles to superconducting circuits to simultaneously exploit the complementary advantages of these two different physical systems (e.g., strong nonlinearity and ease of design of superconducting circuits You and Nori (2005); Schoelkopf and Girvin (2008); Clarke and Wilhelm (2008); You and Nori (2011); Gu et al. (2017); Krantz et al. (2019); Kockum and Nori (2019); Kjaergaard et al. (2020); Blais et al. (2021), and extremely long coherence times of NV spins Bar-Gill et al. (2013)), therefore building an important type of hybrid quantum system Xiang et al. (2013); Clerk et al. (2020). However, the coupling of a single NV spin to a superconducting cavity usually is $g/2\pi\approx 10\,\mathrm{Hz}$ Verdú et al. (2009), which is too weak to allow for a coherent exchange of quantum information between the NV spin and the cavity field. However, the NV spin ensembles typically contains $N\approx 10^{12}$ NV spins, and therefore a collective coupling of $g_{\rm col}/2\pi\approx 10\,\mathrm{MHz}$ can be achieved according to Eq. (72). Such a collective enhancement has been widely demonstrated experimentally Kubo et al. (2010); Amsüss et al. (2011); Kubo et al. (2011, 2012); Putz et al. (2014); Grezes et al. (2014); Astner et al. (2017). The most common setup of such experiments is sketched in Fig. 5(a) Kubo et al. (2010). A diamond single crystal containing a number of NV centers is placed on top of a superconducting coplanar resonator, whose resonance frequency is tunable with an array of four SQUIDs [see Fig. 5(b)]. The collectively enhanced ensemble-cavity coupling can, as seen from Fig. 5(c), cause two clear anticrossings, an indication that the ensemble-cavity system is in the strong coupling regime. Note that here these two anticrossings arise from the fact that there are two distinct microwave transitions between the ground-state spin-triplet sublevels of NV centers, and both can be coupled to the resonator mode. In addition, the collectively enhanced coupling of ensembles of ion spins Schuster et al. (2010); Probst et al. (2013); Wisby et al. (2014) or ${}^{87}$b atoms Hattermann et al. (2017) to a superconducting cavity has also been reported in experiments.
@@
Figure 5: Experimental demonstration of collectively enhanced photon-atom interaction. (a) Photograph of the experimental sample: a nitrogen-vacancy (NV) electronic-spin ensemble is placed on top of a superconducting coplanar waveguide resonator, whose resonance frequency is tuned by modulating the magnetic flux $\Phi$ threading a SQUID array shown in (b). (c) Transmission spectrum $\left|S_{21}\left(\omega\right)\right|$ of the resonator. Two avoided crossings, induced by the strong coupling of the resonator mode to two microwave transitions between the ground-state spin-triplet sublevels of NV centers, can be observed clearly. Here, $B_{\rm NV}$ is an external magnetic field applied to induce the spin Zeeman splitting and $\Phi_{0}=h/2e$ is the flux quantum. Panels (a)-(c) are adapted with permission from Ref. Kubo et al. (2010), Y. Kubo et al., Phys. Rev. Lett. 105, 140502 (2010).
@@
where the first and second terms on the right-hand side describe the dissipative processes of the zero and nonzero momentum modes, respectively. If the coherent dynamics only involves the zero-momentum ($k=0$) mode, then one can focus on only that mode Gelhausen et al. (2017); Shammah et al. (2018); Macrì et al. (2020); that is,
@@
which increases proportionally to the number, $N$, of atoms in the ensemble. It has been shown that even an ensemble weakly coupled to the cavity can induce strong coupling of a single atom to the cavity Schütz et al. (2020).
The collective cooperativity $C_{\rm col}$ in Eq. (78) is only valid for ensembles of independent atoms. When well separated, the atoms can be considered as independent. But when the spacing of these atoms is very small, their dipole-dipole interaction and their collective dissipative dynamics need to be taken into account. In this case, it has been suggested that the collective cooperativity can be further enhanced Plankensteiner et al. (2017, 2019), with an effective collective cooperativity given by
@@ 2.2.3 Other amplification mechanisms
Usually, improving both the quality factor $Q$ and the mode volume $V$ for the same cavity remains challenging, due to the diffraction limit. This means that it is difficult to achieve simultaneously a small cavity decay rate $\kappa$ and a strong photon-atom interaction $g$, which limits the cooperativity $C$. Plasmonic particles (i.e., metal nanoparticles) driven by an external field can produce intense localized fields near them Tame et al. (2013). Many cavities have already utilized such a mechanism to localize an electromagnetic field into a region of nanometer scale and, as a result, to significantly decrease the mode volume of these cavities Seo et al. (2009); Kuttge et al. (2010); Hu and Weiss (2016); Santhosh et al. (2016); Chikkaraddy et al. (2016). This yields an enhancement in the coupling of an atom to the cavity field when the atom is placed closed to the plasmonic particles, and in turn drives the system to the strong coupling regime. For example, for whispering-gallery-mode cavities with an ultrahigh $Q$ but relatively large $V$, the resulting cooperativity is enhanced by approximately two orders of magnitude compared to that obtained in the case of a bare cavity Xiao et al. (2012); Peng et al. (2017). This type of enhancement has been used, e.g., to generate indistinguishable single photons Wein et al. (2018) and quantum entanglement Liu et al. (2013).
@@
In addition, it has been demonstrated that coupling two different cavities, one with a low $Q$ (i.e., with a large cavity loss rate) and the other with a high $Q$ (i.e., with a small cavity loss rate) can effectively realize a high-$Q$ and small-$V$ cavity Liu et al. (2014). The atom is assumed to be coupled to the low-$Q$ cavity. Due to a large detuning and, thus, a low photon occupation, the low-$Q$ cavity mode can be eliminated adiabatically, yielding an effective interaction between the atom and the high-$Q$ cavity mode. This effective atom-cavity system combines the respective advantages of these two cavities (i.e., a high $Q$ and a small $V$). The strong coupling regime, characterized by $C>1$, can then be reached, as long as the detuning of the low-$Q$ cavity is large enough.
@@ 2.3 Amplified Kerr-type light-matter interactions via quadrature squeezing
Phase shifts induced by Kerr-type effects are typically very small when dealing with single photons Venkataraman et al. (2013). A number of experiments have demonstrated the feasibility to generate and observe cross-Kerr phase shifts even on the order of a few tens of degrees per photon, which can enable performing at least limited quantum logic gates. These include experiments based on cavity QED using single atoms Turchette et al. (1995), quantum dots Fushman et al. (2008), or atomic ensembles Beck et al. (2016), circuit QED Hoi et al. (2013), and photonics using optical fibers Matsuda et al. (2009). Anyway, much larger phase shifts at the single-photon level are in high demand in order to harness the full power of the Kerr or Kerr-type effects for quantum computing.
As demonstrated in Ref. Bartkowiak et al. (2014), it is possible, at least theoretically, to boost a cross-Kerr phase shift to an arbitrary value by employing sequentially either one- or two-mode quadrature-squeezing operations. Utilizing such Kerr amplification techniques may prove valuable in implementing quantum nondemolition (QND) measurements Imoto et al. (1985) or practical quantum-optical entangling gates, like a deterministic Fredkin gate Milburn (1989) or a conditional phase (CPHASE) gate Chuang and Yamamoto (1995); Brod and Combes (2016), which rely on giant Kerr nonlinear interactions at the level of individual photons.
There is yet another fundamental advantage of the approach proposed in Ref. Bartkowiak et al. (2014); namely, this method shows the feasibility of enhancing (at least some types of) higher-order nonlinear interactions by applying lower-order nonlinear effects. Indeed, the Kerr effect generated in an nonlinear medium is proportional to its third-order susceptibility $\chi^{(3)}$, while squeezing generation depends on the second-order susceptibility $\chi^{(2)}$.
In this section, we recall the amplification method and circuits proposed in Ref. Bartkowiak et al. (2014). The method was developed based on a vector coherent-state theory and applied physical operations that adhere to the commutation relations associated with the SU(1,1) generators. Thus, let us consider cross-Kerr-type nonlinear interaction between a two-level atom $a$ and an optical mode $b$, as described by the effective Hamiltonian ($\hbar=1$)
@@
Figure 6: Circuits of Ref. Bartkowiak et al. (2014) for amplifying the cross-Kerr interaction between a two-level atom (in path $a$) and optical modes (in paths $b$ and $c$) by sequentially applying quadrature squeezing: (a) single-mode squeezing $S_{1}=S_{b}(\theta_{1})$ and $S_{2}=S_{b}(\theta_{2})$, according to Eq. (83); or (b) two-mode squeezing $S_{1}=S_{bc}(\theta_{1})$ and $S_{2}=S_{bc}(\theta_{2})$, according to Eq. (91), separated with proper phase shifts (PSs). The SWAP gates are represented by the lines connecting the $\times$ symbols.
@@
The single-optical-mode amplification method, described by Eq. (83), was generalized in Ref. Bartkowiak et al. (2014) to the two-optical-mode case. Specifically, the cross-Kerr interaction, given in Eq. (80), is also assumed between the atom $a$ and another optical mode $c$, i.e.,
@@
Figure 7: Experimental setup and circuits of Ref. Burd et al. (2021) for demonstrating parametrically amplified light-matter interactions. (a) Photograph of the surface-electrode ion trap. (b) False-color SEM image of the central region, boxed in (a), of the ion trap. Two ${}^{25}{\rm Mg}^{+}$ ions are trapped $\sim 30\,\mathrm{\mu m}$ above the trap surface. The blue electrodes provide the qubit control, while the red electrodes provide the trapping rf and the parametric modulation for the motional mode. (c) Circuit diagram for parametric modulation of the ion motional frequency. BPF refers to a band-pass filter. Panels (a)-(c) were provided by Dr. Daniel H. Slichter from NIST.
@@ 2.4 Experimental demonstrations of parametrically amplified light-matter interactions
In 2021, an experimental demonstration of parametrically squeezing a bosonic mode to enhance the generation of quantum entanglement between qubits was reported in a trapped-ion system in Ref. Burd et al. (2021) by a group at the National Institute of Standards and Technology (NIST). It has further been experimentally shown, by the same group in Ref. Burd et al. (2023) in 2023, that the use of parametric squeezing can realize the amplification of the system Hamiltonian, even without the precise knowledge of this Hamiltonian Arenz et al. (2020). Recently, the parametrically amplified dispersive interaction between an atom and a squeezed microwave cavity mode was also demonstrated in a superconducting-circuit experiment in Ref. Villiers et al. (2022). Below, we introduce these experiments in more detail.

### v2

@@ 2.2.2 Collective amplification
In Sec. 2.2.1, we showed that a detuned parametric driving of a cavity enables the coupling between an atom and the squeezed cavity mode to be enhanced exponentially. Because exponentially many photons are converted into a single-photon state of the squeezed cavity mode, this parametric enhancement of the atom-field coupling can thus be understood as originating from the coupling of a single atom to many photons.
In this section, we introduce an opposite enhancement mechanism, which is based on the coupling of a single photon to an ensemble containing many atoms or spins. To proceed, let us assume that the ensemble contains $N$ identical two-level atoms. The coupling between the ensemble and a cavity mode is described by the interaction Hamiltonian
@@
Let us now apply the Holstein-Primakoff transformation [299]
$S_{-}=\;\sqrt{N-s^{\dagger}s}s,\quad S_{+}=\;s^{\dagger}\sqrt{N-s^{\dagger}s},\quad{\rm and}\quad S_{z}=-N/2+s^{\dagger}s,$ (69)
where $s$ and $s^{\dagger}$ are the bosonic annihilation and creation operators, which satisfy the commutation relation $\left[s,s^{\dagger}\right]=1$. In the low-excitation regime, where the average number of excited atoms is much smaller than the total number of atoms (i.e., $\langle s^{\dagger}s\rangle\ll N$), the operators $S_{-}$ and $S_{+}$ are further simplified to
$S_{-}\approx\;\sqrt{N}s\quad{\rm and}\quad S_{+}\approx\;\sqrt{N}s^{\dagger},$ (70)
respectively. It is seen that the collective spin behaves as a quantum harmonic oscillator. In this case, the Hamiltonian $H_{\rm ens}^{\rm int}$ becomes
@@
The collective strong coupling between light and matter was first demonstrated experimentally with an ensemble of Rydberg atoms in Ref. [300], and has been a fundamental building block of quantum repeaters for long-distance quantum communication [301, 302, 303, 304, 305, 306, 307, 308, 309, 310, 311, 312]. Moreover, a collective enhancement of light-matter interactions allows one to generate various nonclassical states in large ensembles, e.g., spin-squeezed states [313, 314, 315, 316, 317, 318, 319, 320, 173, 321, 322, 323] and atomic Schrödinger cat states [324, 325, 203]. Thus, it also forms an essential ingredient of quantum metrology for high-precision measurements [326].
@@
Recently, much attention has been focused on coupling nitrogen-vacancy (NV) electronic-spin ensembles to superconducting circuits to simultaneously exploit the complementary advantages of these two different physical systems (e.g., strong nonlinearity and ease of design of superconducting circuits [6, 7, 8, 9, 10, 11, 12, 13, 14], and extremely long coherence times of NV spins [327]), therefore building an important type of hybrid quantum system [328, 329]. However, the coupling of a single NV spin to a superconducting cavity usually is $g/2\pi\approx 10\,\mathrm{Hz}$ [330], which is too weak to allow for a coherent exchange of quantum information between the NV spin and the cavity field. However, the NV spin ensembles typically contains $N\approx 10^{12}$ NV spins, and therefore a collective coupling of $g_{\rm col}/2\pi\approx 10\,\mathrm{MHz}$ can be achieved according to Eq. (72). Such a collective enhancement has been widely demonstrated experimentally [331, 332, 333, 334, 335, 336, 337]. The most common setup of such experiments is sketched in Fig. 5(a) [331]. A diamond single crystal containing a number of NV centers is placed on top of a superconducting coplanar resonator, whose resonance frequency is tunable with an array of four SQUIDs [see Fig. 5(b)]. The collectively enhanced ensemble-cavity coupling can, as seen from Fig. 5(c), cause two clear anticrossings, an indication that the ensemble-cavity system is in the strong coupling regime. Note that here these two anticrossings arise from the fact that there are two distinct microwave transitions between the ground-state spin-triplet sublevels of NV centers, and both can be coupled to the resonator mode. In addition, the collectively enhanced coupling of ensembles of ion spins [338, 339, 340] or 87Rb atoms [341] to a superconducting cavity has also been reported in experiments.
@@
Figure 5: Experimental demonstration of collectively enhanced photon-atom interaction. (a) Photograph of the experimental sample: a nitrogen-vacancy (NV) electronic-spin ensemble is placed on top of a superconducting coplanar waveguide resonator, whose resonance frequency is tuned by modulating the magnetic flux $\Phi$ threading a SQUID array shown in (b). (c) Transmission spectrum $\left|S_{21}\left(\omega\right)\right|$ of the resonator. Two avoided crossings, induced by the strong coupling of the resonator mode to two microwave transitions between the ground-state spin-triplet sublevels of NV centers, can be observed clearly. Here, $B_{\rm NV}$ is an external magnetic field applied to induce the spin Zeeman splitting and $\Phi_{0}=h/2e$ is the flux quantum. Panels (a)-(c) are adapted with permission from Ref. [331], Y. Kubo et al., Phys. Rev. Lett. 105, 140502 (2010).
@@
where the first and second terms on the right-hand side describe the dissipative processes of the zero and nonzero momentum modes, respectively. If the coherent dynamics only involves the zero-momentum ($k=0$) mode, then one can focus on only that mode [342, 343, 173]; that is,
@@
which increases proportionally to the number, $N$, of atoms in the ensemble. It has been shown that even an ensemble weakly coupled to the cavity can induce strong coupling of a single atom to the cavity [344].
The collective cooperativity $C_{\rm col}$ in Eq. (78) is only valid for ensembles of independent atoms. When well separated, the atoms can be considered as independent. But when the spacing of these atoms is very small, their dipole-dipole interaction and their collective dissipative dynamics need to be taken into account. In this case, it has been suggested that the collective cooperativity can be further enhanced [345, 346], with an effective collective cooperativity given by
@@ 2.2.3 Other amplification mechanisms
Usually, improving both the quality factor $Q$ and the mode volume $V$ for the same cavity remains challenging, due to the diffraction limit. This means that it is difficult to achieve simultaneously a small cavity decay rate $\kappa$ and a strong photon-atom interaction $g$, which limits the cooperativity $C$. Plasmonic particles (i.e., metal nanoparticles) driven by an external field can produce intense localized fields near them [92]. Many cavities have already utilized such a mechanism to localize an electromagnetic field into a region of nanometer scale and, as a result, to significantly decrease the mode volume of these cavities [347, 348, 349, 350, 351]. This yields an enhancement in the coupling of an atom to the cavity field when the atom is placed closed to the plasmonic particles, and in turn drives the system to the strong coupling regime. For example, for whispering-gallery-mode cavities with an ultrahigh $Q$ but relatively large $V$, the resulting cooperativity is enhanced by approximately two orders of magnitude compared to that obtained in the case of a bare cavity [352, 353]. This type of enhancement has been used, e.g., to generate indistinguishable single photons [354] and quantum entanglement [355].
@@
In addition, it has been demonstrated that coupling two different cavities, one with a low $Q$ (i.e., with a large cavity loss rate) and the other with a high $Q$ (i.e., with a small cavity loss rate) can effectively realize a high-$Q$ and small-$V$ cavity [356]. The atom is assumed to be coupled to the low-$Q$ cavity. Due to a large detuning and, thus, a low photon occupation, the low-$Q$ cavity mode can be eliminated adiabatically, yielding an effective interaction between the atom and the high-$Q$ cavity mode. This effective atom-cavity system combines the respective advantages of these two cavities (i.e., a high $Q$ and a small $V$). The strong coupling regime, characterized by $C>1$, can then be reached, as long as the detuning of the low-$Q$ cavity is large enough.
@@ 2.3 Amplified Kerr-type light-matter interactions via quadrature squeezing
Phase shifts induced by Kerr-type effects are typically very small when dealing with single photons [357]. A number of experiments have demonstrated the feasibility to generate and observe cross-Kerr phase shifts even on the order of a few tens of degrees per photon, which can enable performing at least limited quantum logic gates. These include experiments based on cavity QED using single atoms [358], quantum dots [359], or atomic ensembles [360], circuit QED [361], and photonics using optical fibers [362]. Anyway, much larger phase shifts at the single-photon level are in high demand in order to harness the full power of the Kerr or Kerr-type effects for quantum computing.
As demonstrated in Ref. [210], it is possible, at least theoretically, to boost a cross-Kerr phase shift to an arbitrary value by employing sequentially either one- or two-mode quadrature-squeezing operations. Utilizing such Kerr amplification techniques may prove valuable in implementing quantum nondemolition (QND) measurements [363] or practical quantum-optical entangling gates, like a deterministic Fredkin gate [364] or a conditional phase (CPHASE) gate [365, 366], which rely on giant Kerr nonlinear interactions at the level of individual photons.
There is yet another fundamental advantage of the approach proposed in Ref. [210]; namely, this method shows the feasibility of enhancing (at least some types of) higher-order nonlinear interactions by applying lower-order nonlinear effects. Indeed, the Kerr effect generated in an nonlinear medium is proportional to its third-order susceptibility $\chi^{(3)}$, while squeezing generation depends on the second-order susceptibility $\chi^{(2)}$.
In this section, we recall the amplification method and circuits proposed in Ref. [210]. The method was developed based on a vector coherent-state theory and applied physical operations that adhere to the commutation relations associated with the SU(1,1) generators. Thus, let us consider cross-Kerr-type nonlinear interaction between a two-level atom $a$ and an optical mode $b$, as described by the effective Hamiltonian ($\hbar=1$)
@@
Figure 6: Circuits of Ref. [210] for amplifying the cross-Kerr interaction between a two-level atom (in path $a$) and optical modes (in paths $b$ and $c$) by sequentially applying quadrature squeezing: (a) single-mode squeezing $S_{1}=S_{b}(\theta_{1})$ and $S_{2}=S_{b}(\theta_{2})$, according to Eq. (83); or (b) two-mode squeezing $S_{1}=S_{bc}(\theta_{1})$ and $S_{2}=S_{bc}(\theta_{2})$, according to Eq. (91), separated with proper phase shifts (PSs). The SWAP gates are represented by the lines connecting the $\times$ symbols.
@@
The single-optical-mode amplification method, described by Eq. (83), was generalized in Ref. [210] to the two-optical-mode case. Specifically, the cross-Kerr interaction, given in Eq. (80), is also assumed between the atom $a$ and another optical mode $c$, i.e.,
@@
Figure 7: Experimental setup and circuits of Ref. [206] for demonstrating parametrically amplified light-matter interactions. (a) Photograph of the surface-electrode ion trap. (b) False-color SEM image of the central region, boxed in (a), of the ion trap. Two ${}^{25}{\rm Mg}^{+}$ ions are trapped $\sim 30\,\mathrm{\mu m}$ above the trap surface. The blue electrodes provide the qubit control, while the red electrodes provide the trapping rf and the parametric modulation for the motional mode. (c) Circuit diagram for parametric modulation of the ion motional frequency. BPF refers to a band-pass filter. Panels (a)-(c) were provided by Dr. Daniel H. Slichter from NIST.
@@ 2.4 Experimental demonstrations of parametrically amplified light-matter interactions
In 2021, an experimental demonstration of parametrically squeezing a bosonic mode to enhance the generation of quantum entanglement between qubits was reported in a trapped-ion system in Ref. [206] by a group at the National Institute of Standards and Technology (NIST). It has further been experimentally shown, by the same group in Ref. [207] in 2023, that the use of parametric squeezing can realize the amplification of the system Hamiltonian, even without the precise knowledge of this Hamiltonian [260]. Recently, the parametrically amplified dispersive interaction between an atom and a squeezed microwave cavity mode was also demonstrated in a superconducting-circuit experiment in Ref. [209]. Below, we introduce these experiments in more detail.